Racal RA-117
Unit N0144 Log

Serial Number:N0144
Year of Manufacture: 1963 ("1963" stamped on 1.7 MHz crystal).
Date of Receipt:2007
State upon Receipt: Non-functional. Several missing or broken component boards. 3 front panel switches gone. Power supply circuitry a mess. Paper capacitors snipped out. Repairs in 40MHz bandpass filter.

2009 May 05: 3/4 mixer subchassis stripped completely and cleaned: 8 hours.
2009 May 06-08: 3/4 mixer reconstructed: 10+4+3+2 hours.
All new RC, except C336A (10pF) and C338 (470pF). R302 replaced with 56 ohms instead of 47. Tested in N0115. Functional. L302 adjusted for null. 1.7MHz tweaked.
2009 Jul 01: VFO-1 rebuild completed.
All new RC, except 3 mica and 2 tubular ceramic caps. RF input coax replaced with piece of RG-? (50 ohm). Oscillator output lead replaced.
2009 Oct 29: Seized screw stub in VFO-2 removed.
2011 Feb 18~: Work started on VFO-2.
2011 Feb 23: VFO-2 rebuild completed.
3 original micas left in use.
2011 Jul: IF strip, BFO reconstructed.
2012 Jan: Crystal calibrator and System switch obtained from surplus sales. Crystal calibrator is actually from an RA-17. Uses Whitworth screws. Crystal calibrator reconstructed.
2012 Feb: Main chassis reassembly started, harmonics and 2nd mixer compartments reconstructed.
2012 Nov: Audio reconstructed, all-new wiring harness, main chassis reconstruction completed. Two poles (one wafer) on SYSTEM switch eliminated by removing switching of B+ to 2nd Mixer (S5.3L) and incorporating modified AVC circuitry which combines the two AVC poles into one. Wired for fixed bias to RF stage (no AVC to RF) for use with 6DJ8 rather 6ES8. New tube set installed, except V14 & V17 previous 5749's, V15 was GE in Sylvania box, V7 new but missing marking.
2012 Dec 05: Powered up. Audio, IF, BFO, crystal calibrator, VFO-2, VFO-1 all operating. VFO-2 alignment tweaked, 6KHz spread. VFO-1 aligned. No reception. 37.5 MHz does not seem to be making it through bandpass filter.
2012 Dec 05: Attempt to align VFO-1 dial to proper stop. Hex head of set-screws for dial hub stripped, unsuccessful attempt to drill out one screw. Stop pin noted to be bent, straightened, may be enough to shift dial sufficiently.
2012 Dec 12: Collection of tube shields assembled, many originals had been missing. Two 7-pin-Shorts used are steel/alloy, others are aluminum. Could still do to obtain 3 9-pin-Medium, 6 7-pin-Short.
2012 Dec 12: New top and bottom covers constructed. Reassembly completed, Philips head screws used for assorted missing screws.
2012 Dec 14: Check bandpass filters by tuning VFO-1 to 40 MHz and down to 37.5 MHz. Signals seem to make it through. Some tweaking of filters done, for better or worse. Feeding 500KHz-2MHz 100mV signals into antenna input do produce carrier response, with Wadley Loop tuned at 0 and 1MHz, so Wadley Loop appears to be working. Suspicion falls on RF front end.
2012 Dec 23: IF and 3/4 mixer checked with 100 KHz signals fed in at stages, in accordance with reference test data in manual. Seems to be about right: < 1mV produces 100uA on meter. Feeding 2.5 MHz at TP3 does not produce expected result, while fed at input to 2-3 MHz bandpass filter produces reasonable result. Circuitry checked around V9: connection for plate load on wrong pin (no function pin 6 instead of plate on pin 7). Corrected. Receiver now well and functional.
2012 Dec 25: VFO-2 aligned. Initial error was +3/-5 KHz, but fairly linear in the error, so no vane-tweaking needed. Frequency spread was wider than it should be though, so base capacitance added: C306 and adjusting C302 shaft relative to drive split gear by loosening clamp screws on drive split gear. Now within +0.3/-0.8 KHz.
2012 Dec 25: While unit is now functioning adequately, fine alignment of bandpass filters and IFs remains to be done.
2015 Jul 15: Weak reception and seemingly intermittent around the MHz tuning. 37.5 MHz BPF examined for levels. In the course of checking adjustments, one section noted to have no effect on output. C45 (33pf red LEMCO mica) faulty, replaced with 5% mica. Levels found are recorded in site alignment section re 37.5MHz BPF.
2015 Jul 15: KHz tuning off by several KHz (pointer must be adjusted to left). Film strip moved one sprocket hole to right, C306 tweaked. Now within ~ 1KHz except at 0 and 1000.

  .. | Assessment | Unit N16 Log | Unit N0115 Log | Unit N0144 Log | Unit N0714 Log | ..
Racal RA-17 & RA-117